OpenStudy (anonymous):

The weights (in pounds) of six dogs are listed below. 17 56 85 38 138 98 The measure of center is found to be 72 lb. Which measure of center is used?

OpenStudy (anonymous):

mean midrange median mode

OpenStudy (anonymous):

@surjithayer

OpenStudy (amistre64):

you might want to simply go with finding the the value of the options to compare with.

OpenStudy (anonymous):

Ohh I see what you mean. I'll try that.

OpenStudy (amistre64):

i would say the mean is the 'hardest' to compute, so work out the other 3, if its not them then it has to be the remaining option

OpenStudy (anonymous):

Ok, thank you :).

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

It is the mean ^_^

OpenStudy (amistre64):

lets try it :) mode is number used most often ... so no mode. median is the number in the middle of the sorted pack: 17 38 56 85 98 138 ^ or average of middle 2: 56+85 = 141, and 72*2 = 144 so that aint it midrange ... how do we compute that one?

OpenStudy (amistre64):

average of high and low 138+17 = 215 so yeah, mean seems fairest

OpenStudy (anonymous):

Yeah. I did the 3 on paper.. However I did mess up on median, but I got the same result lol.

OpenStudy (amistre64):

155 not 215 but stil fails it lol

OpenStudy (anonymous):

Thank you very much amistre :)

OpenStudy (amistre64):

youre welcome